OpenStudy (anonymous):

Solve: 3/x+1/2=5/6 a. 12 b. 9 c. 6 d. 3

hero (hero):

Hint, multiply the middle equation by 3/3

hero (hero):

You'll end up with \(\large\frac{3}{x} + \frac{3}{6} = \frac{5}{6}\)

hero (hero):

Now we can subtract 3/6 from both sides to get: \(\frac{3}{x} = \frac{5}{6} - \frac{3}{6}\)

OpenStudy (anonymous):

9!

hero (hero):

Combining fractions on the right side we get: \(\frac{3}{x} = \frac{2}{6}\)

hero (hero):

Simplifying the right side we get \(\frac{3}{x} = \frac{1}{3}\)

hero (hero):

Cross multiplying we get x = 9

OpenStudy (anonymous):

thanks!